Exercise 1.5. In one type of state lottery 5 distinct numbers are picked fro 1,2,3, 40 uniformly at random. Describe a sample space 2 and a probability measure P to model this experiment. (b) What is the probability that out of the five picked numbers exactly three will be even? Exercise 1.6. We have an urn with 3 green and 4 yellow balls. We ch

For 1.5, I need to do this with and without replacement. Thank you
Answer #1

With replacement:

(a)

All 40 numbers are equally likely and numbers can be repeated so each number out of 5 can be selected in 40 ways. So number of outcomes in sample space will be

40-102400000

Since each outcome is equally likely so probability of each outcome is 1 / 102400000.

(b)

Out of 40 numbers 20 are even and 20 are odd so number of ways of selecting 3 even numbers and 2 odd numbers is

(20*20*20)*(20*20) = 3200000

Therefore the probability of selecting 3 even numbers is

P(even numbers) = 3200000 / 102400000 = 0.03125

-------------------------------------------

Without replacement:

(a)

All 40 numbers are equally likely and numbers cannot be repeated so number of outcomes in sample space will be

C(40, 5)--一40! 51(40-5),-658008

Since each outcome is equally likely so probability of each outcome is 1 / 658008.

(b)

Out of 40 numbers 20 are even and 20 are odd so number of ways of selecting 3 even numbers and 2 odd numbers is

20! C(20,3)C(20,2) 20! 3!( 20-31 21(20-2),一216600

Therefore the probability of selecting 3 even numbers is

P(even numbers) = 216600 / 658008= 0.3292